# Joint Scheduling of DER under Demand Charges: Structure and Approximation

### v1

@@ IV-A Relaxation of the DER Co-Optimization Problem
We consider a large capacity regime as a relaxation of the battery capacity constraint (9e). Observing that $e_{t}$ and $\mathbf{d}_{t}$ consistently appear together in the expression $\mathbf{1}^{\top}\mathbf{d}_{t}+e_{t}$, we introduce a helper function $h_{t}(v_{t})$:
$h_{t}(v_{t}):=\max\{U_{t}(\bm{d}_{t})+\gamma e_{t}:\bm{1}^{\top}\bm{d}_{t}+e_{t}=v_{t},\bm{0}\preceq\bm{d}_{t}\preceq\overline{\bm{d}},-\underline{e}\leq e_{t}\leq\overline{e}\},$ (15)
where $v_{t}$ represents the net consumption excluding renewables at time step $t$. The function $h_{t}(v_{t})$ is differentiable, monotonically increasing, and concave for all time steps $t$. For any $v_{t}\in[-\underline{e},\overline{e}+\bm{1}^{\top}\bm{\overline{d}}]$, there exists a valid pair $(e_{t},\mathbf{d}_{t})$ that achieves $h_{t}(v_{t})$ with $\mathbf{1}^{\top}\mathbf{d}_{t}+e_{t}=v_{t}$. This equivalence allows us to optimize over $v_{t}$ after relaxing constraint (9e).
Using the helper function (15), we formulate the relaxed problem as:
$\max_{v_{0},\ldots,v_{T-1}}\quad$ $\sum_{t=0}^{T-1}h_{t}(v_{t})-p_{t}^{+}[v_{t}-g_{t}]^{+}+p_{t}^{-}[v_{t}-g_{t}]^{-}-p[v_{t}-g_{t}-c_{t}]^{+}$
Subject to $-\underline{e}\leq v_{t}\leq\overline{e}+\bm{1}^{\top}\overline{\bm{d}}$ (16a)
$c_{t+1}=\max\{c_{t},v_{t}-g_{t}\},$ (16b)
To eliminate the temporal dependencies introduced by constraint (9h), we consider a fixed peak demand bound $c$ in the large capacity regime (16):
@@
This reformulation replaces the recursive peak demand update (16b) with a single-step upper-bound constraint (17c), effectively removing temporal dependencies while preserving the optimal policy structure. We show that the optimal policy of (16) is preserved under this relaxation.
@@
By Thm. 2, we can compute the optimal policy for (16) by searching for the optimal $c^{*}$ and solving (17).
IV-B Solution under Given Peak Demand
@@
$v_{t}^{\ast}(g_{t})=\begin{cases}-\underline{e},\quad\quad\quad\quad\quad\ v^{\dagger}(g_{t})\leq-\underline{e},\\ v_{t}^{\dagger}(g_{t}),\quad\quad\quad\quad-\underline{e}\leq v^{\dagger}(g_{t})\leq\min\{c+g_{t},\overline{e}+\bm{1}^{\top}\overline{\bm{d}}\},\\ \min\{c+g_{t},\overline{e}+\bm{1}^{\top}\overline{\bm{d}}\},\min\{c+g_{t},\overline{e}+\bm{1}^{\top}\overline{\bm{d}}\}\leq v^{\dagger}(g_{t}),\end{cases}$ (18)
where the unconstrained optimum $v_{t}^{\dagger}(g_{t})$ is the solution without constraints (17b) and (17c), given by
$v_{t}^{\dagger}(g_{t})=\begin{cases}(h^{\prime})^{-1}(p_{t}^{-}),&(h^{\prime})^{-1}(p_{t}^{-})<g_{t},\\ g_{t},&(h^{\prime})^{-1}(p_{t}^{+})\leq g_{t}\leq(h^{\prime})^{-1}(p_{t}^{-}),\\ (h^{\prime})^{-1}(p_{t}^{+}),&g_{t}<(h^{\prime})^{-1}(p_{t}^{+}).\end{cases}$ (19)
@@
With fixed peak demand, the optimization decomposes into independent subproblems at each time step $t=0,1,\cdots,T-1$:
@@
$v_{t}^{\dagger}(g_{t})=\arg\max_{v}~{}~{}h_{t}(v)-p_{t}^{+}[v-g_{t}]^{+}+p_{t}^{-}[v-g_{t}]^{-}=\begin{cases}(h_{t}^{\prime})^{-1}(p_{t}^{-}),&(h_{t}^{\prime})^{-1}(p_{t}^{-})<g_{t},\\ g_{t},&(h_{t}^{\prime})^{-1}(p_{t}^{+})<g_{t}<(h_{t}^{\prime})^{-1}(p_{t}^{-}),\\ (h_{t}^{\prime})^{-1}(p_{t}^{+}),&g_{t}<(h_{t}^{\prime})^{-1}(p_{t}^{+}).\end{cases}$
@@
The optimal policy has a clear economic interpretation based on the relationship between renewable generation and consumption needs. The unconstrained solution $v_{t}^{\dagger}(g_{t})$ balances the marginal utility of consumption against energy trading costs. When $(h_{t}^{\prime})^{-1}(p_{t}^{-})<g_{t}$, the renewable generation $g_{t}$ exceeds the energy demanded by consuming at the selling price $\pi^{-}$ and charging the battery. In this case, the optimal consumption $v_{t}^{\dagger}(g_{t})=(h_{t}^{\prime})^{-1}(p_{t}^{-})$ is set such that the marginal utility equals the selling price, meaning excess generation beyond this consumption level should be sold to the grid. When $(h_{t}^{\prime})^{-1}(p_{t}^{+})\leq g_{t}\leq(h_{t}^{\prime})^{-1}(p_{t}^{-})$, the renewable generation exactly matches the optimal consumption level, so $v_{t}^{\dagger}(g_{t})=g_{t}$ and no energy trading occurs. When $g_{t}<(h_{t}^{\prime})^{-1}(p_{t}^{+})$, renewable generation is insufficient relative to the consumption level that would justify purchasing energy at price $p_{t}^{+}$. Here, the optimal consumption $v_{t}^{\dagger}(g_{t})=(h_{t}^{\prime})^{-1}(p_{t}^{+})$ sets marginal utility equal to the purchase price, requiring additional energy to be bought from the grid. The final policy $v_{t}^{*}(g_{t})$ then projects this unconstrained optimum onto the feasible region, ensuring that consumption respects both battery operation limits and peak demand constraints.
IV-C Algorithmic Search of the Optimal Peak Demand
@@ IV-D Large Storage Peak Searching (LSPS) Algorithm
The LSPS algorithm, in Alg. 1, efficiently approximates the problem (9) by combining the relaxation techniques from the previous sections. It leverages the key structural properties established in the previous sections: the decomposition enabled by the large capacity relaxation, the analytical solution for fixed peak demand, and the concavity of the objective function in the peak demand parameter. By searching over the discrete candidate set $\mathcal{C}$, we avoid solving the full dynamic optimization problem and rather solve a linear complexity algorithm, as formalized below, while maintaining a small solution gap compared to the optimal.
@@
6:Project $v_{t}^{*}$ back to variables $(e_{t}^{*},\mathbf{d}_{t}^{*})$ using (15)
@@
Alg. 1 has computational complexity $\mathcal{O}(T)$.
@@
Empirical analysis on the performance of the LSPS algorithm is presented in Sec. V.
@@ V Case Study
Using two rich datasets, and considering a one-day operation horizon $T=24$, we evaluate the DER co-optimization performance of the proposed LSPS algorithm and an RL controller using a Proximal Policy Optimization (PPO) algorithm compared to a theoretical maximum that assumes perfect knowledge of future renewables solving Eq. 9. The performance is measured by the solution gap, which is the difference between the cumulate reward of the algorithms and the theoretical upper bound. Experiments were run on a server with 1 AMD Ryzen Threadripper 3990X 64-Core Processor and 4 Nvidia RTX A4000 GPUs.
@@ V-A Dataset Description
(a) Daily net demand for HouseZero (Jun 2022 – May 2024)
(b) Daily net demand for BDGP building (Jan 2016 – Dec 2017)
@@
Two open-source datasets from Harvard HouseZero [24] and Building Data Genome Project (BDGP) [25] were used to evaluate the performance of DER scheduling under different algorithms. The two buildings were chosen with distinct characteristics while both are equipped with solar PV. Renewable generation and demand data, resampled to hourly granularity, are used for training and testing the performance of different control strategies. Fig. 2 depicts the net demand distribution for HouseZero and BDGP buildings. The two buildings differ significantly in their energy pattern: HouseZero is a low-energy, ultra-efficient building with a peak demand below 2 kW and a frequently negative net consumption. In contrast, the BDGP building has substantially higher consumption, with net consumption peaking at over 160 kW.
@@ V-B Simulation Setup
We adopt the concave utility function from [26]: $U_{t}(d_{t})=\alpha_{t}d_{t}-\frac{1}{2}\beta_{t}d_{t}^{2}$, where $\alpha_{t}$ and $\beta_{t}$are time-varying and learned apriori using historical energy prices, demands, and demand elasticity [27]. The elasticity was set to -0.1 [28]. To get more crisp insights, we assumed time-invariant NEM parameters, with a $p^{+}=\$0.12kWh$ retail (import) rate and a $p^{-}=\$0.06/kWh$ sell (export) rate. The daily peak demand charge was set to $p=\$10/kW$. The final salvage value for the remaining energy in the battery was set at $\gamma=\$0.09/kWh$. The baseline battery parameters were selected based on the demand data and size of the building. For HouseZero, $B=5$ kWh and $\underline{e}=\overline{e}=1$ kW were used; while for BDGP building, $B=350$ kWh and $\underline{e}=\overline{e}=50$ kW were used. In both cases, the battery charging/discharging efficiencies were set at $\tau=\rho=0.95$.
@@ V-C Control Strategies
We present five control strategies to schedule DER here:
@@
The backup mode prioritizes using renewables to charge the battery to maintain a high SoC level, which is usually set by the user. The battery discharges only during grid outages, equivalent to the case without a battery system [4] during normal operations.
@@
Mathematically, the storage control under RATP is
$e_{t}(g_{t})=\begin{cases}\max\{-\underline{e},\rho s_{t},-\tilde{d}_{t}\},&\bar{d}_{t}-g_{t}>0\\ \min\{\overline{e},(B-s_{t})/\tau,-\tilde{d}_{t}\}&\bar{d}_{t}-g_{t}\leq 0,\end{cases}$
and $d_{t}=\bar{d}_{t}$ in both cases. We use RATP and Threshold interchangeably.
@@
We use a popular model-free RL method of proximal policy optimization (PPO) algorithms to directly learn the optimal policy for controlling the DER system described. PPO uses an actor-network and a critic network with a multilayer perceptron (MLP) structure. For training, each of the actor and critic networks consists of two hidden layers with 256 neurons per layer. The actor-network predicts the optimal action, while the critic network estimates the optimal value function. We train the networks with a learning rate of $1e^{-4}$, batch size of 64, and 16 epochs per update. Other training parameters are kept as default provided by Stable-baselines3 [29].
@@
LSPS algorithm, described in Alg. 1, offers an efficient and scalable method to co-optimize DER. Following Alg. 1, we first calculate the optimal peak, then, by solving Eq. 17 with the optimal peak, we get the battery and demand actions at each time step.
@@
The theoretical upper bound assumes perfect knowledge of the renewables for the entire scheduling period and solves the constrained convex program in Eq. 9.
@@ V-D Simulation Results
We compare the performance of the control strategies (a)-(d) above to the theoretical optimal algorithm in (e) by varying the following parameters:
@@
  • Export rate of electricity ($p^{-}$): \$$\$0/\text{kWh}$\$$\$0.12/\text{kWh}$ from no to high sell back incentive with $p^{-}=p^{+}$.
  • Peak demand price of electricity ($p$): \$$\$0/\text{kW}$\$$\$10/\text{kW}$ from no penalty on a high peak to a reasonably high peak charge commonly practiced in electricity tariffs.
Tab. III summarizes the overall performance comparisons in terms of the solution gap to the optimal solutions. [^3] Table IV in the Appendix details all cases for reference. Overall, the LSPS algorithm performed closest to the optimal solutions over all scenarios with an average gap of 13.77%, followed by RL of 20.21%, threshold algorithm of 41.93%, and backup mode of 43.75%. The LSPS algorithm was closest to the optimal 34 out of 42 scenarios (81%) while RL was closest to optimal in the remaining 8 scenarios (19%). Furthermore, for almost all cases, control strategies/algorithms showed better performance for the smaller building of HouseZero, with LSPS performing only 4.52% below the theoretical optimal. The higher peak demand and energy consumption presented greater challenges for optimal operations of the DER system.
TABLE III: Summary of different control strategies under test cases
@@
Backup Threshold RL LSPS
@@
Illustrated in Fig. 3, for the first group of test cases varying battery capacity sees the LSPS algorithm performed closest to the theoretical upper bound for both HouseZero and BDGP in Fig. 3(a) and Fig. 3(e). For the HouseZero case with low loads and large battery capacities, the peak was completely removed (0kW), leading the LSPS algorithm to perform as well as the theoretical optimal. Large battery capacity also significantly improved the threshold algorithm performance, even though it significantly fell behind RL and LSPS algorithms when the battery size was more limited. On the other hand, RL followed a similar trend as the LSPS algorithm while performing around 10% worse over all cases. RL was still stable and more adaptive to changes in the environment than the threshold and backup algorithms. For the much larger BDGP building, both RL and LSPS algorithms improved overall with the increased battery capacity while backup and threshold failed to show the expected improvement due to their inflexibility in considering the high peak and energy use in this building.
@@
Second, varying salvage value rate again showed the superior performance of the LSPS algorithm for both HouseZero and BDGP in Fig. 3(b) and Fig. 3(f). The threshold algorithm tended to use up the available battery charge, leading to low salvage value at the end of the day and causing a larger gap to optimal with increased salvage value rate. Meanwhile, for extremely high salvage value cases, the backup case got closer to optimal as the strategy kept the battery full at the end of the day. RL performed relatively well and was closer to the LSPS algorithm and optimal for the relatively smaller HouseZero building. Overall, for both HouseZero and BDGP, the performances exhibited a similar trend across the four control strategies/algorithms.
@@
Third, varying export rates highlighted the difference between HouseZero and BDGP buildings illustrated in Fig. 3(c) and Fig. 3(g). For the BDGP building, the higher energy demand meant the building would never be in net producing regime, leading to flat lines for the control strategies/algorithms. With RL, there were slight variations in performance from the stochasticity during training. For HouseZero, all control strategies/algorithms improved with a higher export rate since it generated more revenue for the building to sell back excessive energy production from solar. LSPS and RL algorithms performed relatively well and were stable across different export rates over the two test buildings.
@@
Lastly, varying peak demand charges saw all control strategies/algorithms change in a similar trend shown in Fig. 3(d) and Fig. 3(h). It was observed that for low peak charges, all control strategies/algorithms performed well, highlighting the complexity of peak demand charges added to optimally managing such a DER system. With higher peak demand charges, all control strategies/algorithms diverged from optimal, with LSPS and RL algorithms performing the best across different scenarios. Additionally, both of them showed a smaller solution gap for the smaller building of HouseZero, indicating the challenges for buildings with higher demand and energy consumption.
@@
Overall, both LSPS and RL showed great robustness and performance across different scenarios with varying battery capacities, salvage values, and prices of electricity. Specifically, LSPS optimization outperformed backup and threshold algorithms by 68.47% and 68.9% while RL outperformed them by 52.49% and 48.01% respectively. LSPS and RL algorithms trended in sync in their gap to optimality, demonstrating their ability to capture system dynamics while the LSPS algorithm is superior in around 81% instances. Moreover, LSPS was optimal with a large battery capacity and close to optimal for low export rate and peak demand charge. With 0 or low peak demand charges, all control strategies/algorithms performed close to optimal. In comparison, the threshold algorithm and backup mode were sensitive to changes in battery capacity, salvage rate, peak demand charge, and export rate when the building entered the net-producing regime. For cases where the battery capacity was large or the peak price was low, the threshold algorithm outperformed RL and approaches optimal with deterministic predictions. However, with a high salvage value and demand price, the threshold algorithm’s performance degraded significantly and became even worse than the backup mode. Backup mode varied mostly linearly in its gap to the optimal case. It performed well only in special cases with high salvage value rates and low peak demand cases.

### v2

@@ IV-A Relaxation of the DER Co-Optimization Problem
We consider a large storage capacity regime as a relaxation of the battery capacity constraint (10e). Observing that $e_{t}$ and $\bm{d}_{t}$ consistently appear together in the expression $\mathbf{1}^{\top}\bm{d}_{t}+e_{t}$, we introduce a helper function $h_{t}(v_{t})$:
$h_{t}(v_{t})=\max\{$ $U_{t}(\bm{d}_{t})+\gamma e_{t}:\bm{1}^{\top}\bm{d}_{t}+e_{t}=v_{t},$ (16)
$\bm{0}\preceq\bm{d}_{t}\preceq\overline{\bm{d}},-\underline{e}\leq e_{t}\leq\overline{e}\},$
where $v_{t}$ represents the net consumption excluding renewables at time step $t$. The function $h_{t}(v_{t})$ is differentiable, monotonically increasing, and concave for all time steps $t$. For any $v_{t}\in[-\underline{e},\overline{e}+\bm{1}^{\top}\bm{\overline{d}}]$, there exists a valid pair $(e_{t},\bm{d}_{t})$ that achieves $h_{t}(v_{t})$ with $\mathbf{1}^{\top}\bm{d}_{t}+e_{t}=v_{t}$. This equivalence allows us to optimize over $v_{t}$ after relaxing constraint (10e).
Using the helper function (16), we formulate the relaxed problem as
$\max_{v_{0},\ldots,v_{T-1}}\quad$ $\sum_{t=0}^{T-1}h_{t}(v_{t})-p_{t}^{+}[v_{t}-g_{t}]^{+}+p_{t}^{-}[v_{t}-g_{t}]^{-}$
$-p[v_{t}-g_{t}-c_{t}]^{+}$ (17a)
Subject to $-\underline{e}\leq v_{t}\leq\overline{e}+\bm{1}^{\top}\overline{\bm{d}}$ (17b)
$c_{t+1}=\max\{c_{t},v_{t}-g_{t}\}.$ (17c)
To eliminate the temporal dependencies introduced by constraint (10h), we consider a fixed peak demand bound $c$ in the large capacity regime (17)
@@
This reformulation replaces the recursive peak demand update (17c) with a single-step upper-bound constraint (18c), effectively removing temporal dependencies while preserving the optimal policy structure. We show that the optimal policy of (17) is preserved under this relaxation.
@@
By Theorem 2, we can compute the optimal policy for (17) by searching for the optimal $c^{*}$ and solving (18).
IV-B Solution for a Given Peak Demand
@@
$v_{t}^{\ast}(g_{t})=$
$\begin{cases}-\underline{e},&\text{if}\quad v^{\dagger}_{t}(g_{t})<-\underline{e},\\ v_{t}^{\dagger}(g_{t}),&\text{if}\quad-\underline{e}\leq v^{\dagger}_{t}(g_{t})\leq\Gamma_{t},\\ \Gamma_{t},&\text{if}\quad\Gamma_{t}<v^{\dagger}_{t}(g_{t}),\end{cases}$ (19)
where $\Gamma_{t}=\min\{c+g_{t},\overline{e}+\bm{1}^{\top}\overline{\bm{d}}\}$ is the upper bound of $v_{t}^{*}(g_{t})$, the unconstrained optimum $v_{t}^{\dagger}(g_{t})$ is the solution without constraints (18b) and (18c), given by
$v_{t}^{\dagger}(g_{t})$
$=$ $\begin{cases}(h^{\prime})^{-1}(p_{t}^{-}),&\text{if}\quad(h^{\prime})^{-1}(p_{t}^{-})<g_{t},\\ g_{t},&\text{if}\quad(h^{\prime})^{-1}(p_{t}^{+})\leq g_{t}\leq(h^{\prime})^{-1}(p_{t}^{-}),\\ (h^{\prime})^{-1}(p_{t}^{+}),&\text{if}\quad g_{t}<(h^{\prime})^{-1}(p_{t}^{+}).\end{cases}$ (20)
@@
With fixed peak demand, the optimization decomposes into independent subproblems at each time step $t=0,1,\ldots,T-1$:
@@
$v_{t}^{\dagger}(g_{t})=\arg\max_{v}h_{t}(v)-p_{t}^{+}[v-g_{t}]^{+}+p_{t}^{-}[v-g_{t}]^{-}$
[…]
@@
The optimal policy has an intuitive economic interpretation based on the relationship between renewable generation and consumption needs. The unconstrained optimal solution $v_{t}^{\dagger}(g_{t})$ balances the marginal utility of consumption against energy costs. When $(h_{t}^{\prime})^{-1}(p_{t}^{-})<g_{t}$, the renewable generation $g_{t}$ exceeds the energy demanded by consuming at the export price $p^{-}$ and charging the battery. In this case, the optimal consumption $v_{t}^{\dagger}(g_{t})=(h_{t}^{\prime})^{-1}(p_{t}^{-})$ is set such that the marginal utility equals the export price, meaning excess generation beyond this consumption level should be sold to the grid. When $(h_{t}^{\prime})^{-1}(p_{t}^{+})\leq g_{t}\leq(h_{t}^{\prime})^{-1}(p_{t}^{-})$, the renewable generation exactly matches the optimal consumption level, so $v_{t}^{\dagger}(g_{t})=g_{t}$ and no energy transactions occur. When $g_{t}<(h_{t}^{\prime})^{-1}(p_{t}^{+})$, renewable generation is insufficient relative to the consumption level that would justify purchasing energy at price $p_{t}^{+}$. Here, the optimal consumption $v_{t}^{\dagger}(g_{t})=(h_{t}^{\prime})^{-1}(p_{t}^{+})$ sets marginal utility equal to the purchase price, requiring additional energy to be bought from the grid. The final policy $v_{t}^{*}(g_{t})$ then projects this unconstrained optimum onto the feasible region, ensuring that consumption respects both battery operation limits and peak demand constraints.
IV-C Algorithmic Search for the Optimal Peak Demand
@@ IV-D Large Storage Peak Searching (LSPS) Algorithm
The LSPS algorithm (Algorithm 1) is designed to solve the scheduling problem offline. For a given scheduling horizon (e.g., 24 hours), the algorithm uses forecasts of exogenous variables, like renewable generation $g_{t}$, for the entire period to solve for a single optimal peak demand, $c^{*}$. This single $c^{*}$ is then used to determine the optimal control actions for each time step within the horizon. The algorithm efficiently approximates the problem (10) by combining the relaxation techniques from the previous sections. It leverages the key structural properties established in the previous sections: the decomposition enabled by the large storage capacity relaxation, the analytical solution for fixed peak demand, and the concavity of the objective function in the peak demand parameter. By searching over the discrete candidate set $\mathcal{C}$, we avoid solving the full dynamic optimization problem and rather solve a linear complexity algorithm, as formalized below, while maintaining a small solution gap compared to the optimal.
@@
6:Project $v_{t}^{*}$ back to variables $(e_{t}^{*},\bm{d}_{t}^{*})$ using (16)
@@
Algorithm 1 has computational complexity $\mathcal{O}(T)$.
@@
Empirical analysis on the performance of the LSPS algorithm is presented in §V. A comparison between the computational cost of the LSPS algorithm and a deterministic optimization that solves (10) offline in one-shot is presented in Appendix -C.
@@ V Case Study
Using two rich datasets, and considering a one-day operation horizon with hourly decisions, $T=24$ hours, we evaluate the DER co-optimization performance of the proposed LSPS algorithm and several benchmarks compared to a theoretical maximum that assumes perfect knowledge of future renewables solving (10). Performance is measured by the solution gap, which is the difference between the cumulative reward of the algorithms and the theoretical upper bound. The experiments were run on a server with 1 AMD Ryzen Threadripper 3990X 64-Core Processor and 4 Nvidia RTX A4000 GPUs.
@@ V-A Dataset Description
(a) Daily net demand data for HouseZero from Jun 2022 to May 2024
(b) Daily net demand data for BDGP building from Jan 2016 to Dec 2017
@@
Two open-source datasets from Harvard HouseZero [30] and Building Data Genome Project (BDGP) [31] were used to evaluate the performance of DER scheduling under different algorithms. The two buildings were chosen with distinct characteristics, while both are equipped with solar PV. Renewable generation and demand data, resampled to hourly granularity, were used to train and test the performance of different control strategies. Fig. 2 shows the net demand distribution for the HouseZero and BDGP buildings. The two buildings differ significantly in their energy pattern: HouseZero is a low-energy, ultra-efficient building with a peak demand below 2 kW and a frequently negative net consumption. In contrast, the BDGP building has substantially higher consumption, with net consumption peaking at over 160 kW.
@@ V-B Simulation Setup
We adopt the concave utility function from [32]: $U_{t}(d_{t})=\alpha_{t}d_{t}-\frac{1}{2}\beta_{t}d_{t}^{2}$, where $\alpha_{t}$ and $\beta_{t}$are time-varying and learned a priori using historical energy prices, demands, and demand elasticity [26]. The price elasticity of consumption was set to 0.1-0.1 [33]. To get more crisp insights, we assumed time-invariant NEM tariff parameters, with a retail (import) rate of $p^{+}=\mathdollar 0.12$/kWh and a sell (export) rate of $p^{-}=\mathdollar 0.06$/kWh. The daily peak demand charge was set to $p=\mathdollar 10$/kW. The final salvage value for the remaining energy in the battery was set at $\gamma=\mathdollar 0.09$/kWh. The baseline battery parameters were selected based on the demand data and the size of the building. For HouseZero, $B=5$ kWh and $\underline{e}=\overline{e}=1$ kW were used; while for BDGP building, $B=350$ kWh and $\underline{e}=\overline{e}=50$ kW were used. In both cases, the battery charging/discharging efficiencies were set at $\tau=\rho=0.95$.
@@ V-C Control Strategies
We consider and compare five distinct control strategies for scheduling DER. The description of each strategy is provided below, while a detailed comparison of their respective advantages and limitations is presented in Appendix -D.
@@
The backup mode prioritizes using renewables to charge the battery to maintain a high SoC level, which is usually set by the user. The battery discharges only during grid outages, equivalent to the case without a battery system [4] during normal operations. Consequently, this control strategy cannot exploit the full potential of jointly optimizing demand and storage.
@@
Mathematically, the storage control under RATP, for a given demand $d_{t}=\hat{d}_{t}$, is
$e_{t}(g_{t})=\begin{cases}\max\{-\underline{e},-\rho s_{t},-\tilde{d}_{t}\},&\tilde{d}_{t}(g)>0\\ \min\{\overline{e},(B-s_{t})/\tau,-\tilde{d}_{t}\}&\tilde{d}_{t}(g)\leq 0,\end{cases}$
for $t=0,1,\ldots,T-1$.
@@
We use a popular model-free RL method of proximal policy optimization (PPO) algorithms to directly learn the optimal policy for controlling the DER system described. PPO uses an actor-network and a critic network with a multilayer perceptron (MLP) structure. For training, each of the actor and critic networks consists of two hidden layers with 256 neurons per layer. The actor-network predicts the optimal action, while the critic network estimates the optimal value function. We train the networks with a learning rate of $1e^{-4}$, batch size of 64, and 16 epochs per update. Other training parameters are kept as the default provided by Stable-baselines3 [34]. Training histories for RL are provided in Appendix -E, showing the convergence of the trained policy that is used in §V-D.
@@
LSPS algorithm, described in Algorithm 1, offers an efficient and scalable method to co-optimize DER. Following Algorithm 1, we first calculate the optimal peak, then, by using Lemma 4 with the optimal peak, we compute the co-optimized battery and demand actions at each time step.
@@
The theoretical upper bound assumes perfect knowledge of the renewables for the entire scheduling period and deterministically solves the constrained convex program in (10) in one-shot while relaxing the simultaneous charging-discharging constraint, giving it a slight advantage over LSPS.
During the prediction stage, RL takes in the current state values, including battery SoC, renewable generation, peak demand, and outputs the optimal actions based on the learned policy. While the LSPS algorithm during prediction takes in the optimal peak demand in addition to the current demand and generation. The optimal actions are generated by solving an optimization problem. During the training stage, a similar level of computation is required as RL learns the system dynamics, including the patterns of the renewable profile with two years of hour-level training data, while LSPS requires models for renewable predictions.
@@ V-D Simulation Results
We compare the performance of the control strategies (a)-(d) above to the theoretical upper bound in (e) by varying the following parameters:
@@
  • Export rate of electricity ($p^{-}$): $\mathdollar 0/\text{kWh}$$\mathdollar 0.12/\text{kWh}$ from no to high sell back incentive with $p^{-}=p^{+}$.
  • Peak demand price of electricity ($p$): $\mathdollar 0/\text{kW}$$\mathdollar 10/\text{kW}$ from no penalty on a high peak to a reasonably high peak charge commonly practiced in electricity tariffs.
Table III summarizes the overall performance comparisons in terms of the solution gap to the optimal solutions. [^5] Table V in Appendix -B provides detailed results for all cases. Overall, the LSPS algorithm performed closest to the optimal solutions over all scenarios with an average gap of 12.86%, followed by RL of 20.21%, RATP algorithm of 41.93%, and backup mode of 43.75%. The LSPS algorithm was closest to the optimal 34 out of 42 scenarios (81%) while RL was closest to optimal in the remaining 8 scenarios (19%). Furthermore, for almost all cases, control algorithms showed better performance for the smaller building of HouseZero, with LSPS performing only 4.52% below the theoretical optimal.
The performance gap between the HouseZero and BDGP buildings is mainly attributed to differences in their peak demand, total energy consumption, and net load profiles. Specifically, the shape of the net load depends on whether the building transitions between net-producing and net-consuming regimes, the timing of these transitions, and the duration of each regime. In the case of HouseZero, lower demand and higher on-site generation result in negative net load during daytime hours, with peak demand occurring in the early morning or late evening. In contrast, the substantially higher demand of the BDGP building prevents it from reaching a net-producing state, leading to peak demand periods during daytime hours.
TABLE III: Summary of solution gap (%) of different control strategies
@@
Backup RATP RL LSPS
@@
Illustrated in Fig. 3, for the first group of test cases varying battery capacity, the LSPS algorithm performed closest to the theoretical upper bound for both HouseZero and BDGP in Fig. 3(a) and Fig. 3(e). For the HouseZero case with low loads and large battery capacities, the peak was completely removed (0 kW), leading the LSPS algorithm to perform as well as the theoretical optimal. Large battery capacity also significantly improved the RATP algorithm performance, even though it significantly fell behind the RL and LSPS algorithms when the battery size was more limited. On the other hand, RL followed a similar trend as the LSPS algorithm while performing around 10% worse over all cases. RL was still stable and more adaptive to changes in the environment than the RATP and backup algorithms. For the much larger BDGP building, both RL and LSPS algorithms improved overall with the increased battery capacity, while backup and RATP failed to show the expected improvement due to their inflexibility in considering the high peak and energy use in this building.
@@
Second, varying salvage value rate again showed the superior performance of the LSPS algorithm for both HouseZero and BDGP in Fig. 3(b) and Fig. 3(f). The RATP algorithm tended to use up the available battery charge, leading to low salvage value at the end of the day and causing a larger gap to optimal with an increased salvage value rate. Meanwhile, for extremely high salvage value cases, the backup case got closer to optimal as the strategy kept the battery full at the end of the day. RL performed relatively well and was closer to the LSPS algorithm and optimal for the relatively smaller HouseZero building. Overall, for both HouseZero and BDGP, the performances exhibited a similar trend across the four control algorithms.
@@
Third, varying export rates highlighted the difference between HouseZero and BDGP buildings, illustrated in Fig. 3(c) and Fig. 3(g). For the BDGP building, the higher energy demand meant the building would never be in a net producing regime, leading to flat lines for the control algorithms. With RL, there were slight variations in performance due to the stochasticity during training. For HouseZero, all control algorithms improved with a higher export rate since it generated more revenue for the building to sell back the excess energy production from solar. LSPS and RL algorithms performed relatively well and were stable across different export rates over the two test buildings.
@@
Lastly, varying peak demand charges saw all control algorithms change in a similar trend shown in Fig. 3(d) and Fig. 3(h). It was observed that for low peak charges, all control algorithms performed well, highlighting the complexity of peak demand charges added to optimally managing such a DER system. With higher peak demand charges, all control algorithms diverged from optimal, with LSPS and RL algorithms performing the best across different scenarios. Additionally, both of them showed a smaller solution gap for the smaller building of HouseZero, indicating the challenges for buildings with higher demand and energy consumption.
In addition to analyzing the sensitivity of the above-mentioned parameters, we compared the performance of the RL and LSPS algorithms under different load and renewable patterns in Tab. IV. Results show that higher generation and lower demand provided higher surplus and a smaller gap to optimal from both RL and LSPS. On average, RL performed with a gap of 17.93% while LSPS performed with a gap of 4.43%. The results demonstrated the effectiveness of the RL and LSPS across various load patterns, highlighting the potential for implementing the algorithm in future long-term real-world tests.
TABLE IV: RL and LSPS performance under different generation and demand scenarios.
Surplus Gap to Optimal
Generation Demand RL LSPS Optimal RL LSPS
75th Percentile 25th Percentile 16.4 18.9 20.3 19.2% 6.9%
50th Percentile 50th Percentile 18.2 21.0 21.9 16.9% 4.1%
25th Percentile 75th Percentile 18.1 21.5 22.0 17.7% 2.3%
@@
Overall, both LSPS and RL showed great robustness and performance across different scenarios with varying battery capacities, salvage values, and prices of electricity. Specifically, LSPS optimization outperformed backup and RATP algorithms by 68.47% and 68.9%, while RL outperformed them by 52.49% and 48.01%, respectively. LSPS and RL algorithms trended in sync in their gap to optimality, demonstrating their ability to capture system dynamics, while the LSPS algorithm is superior in around 81% instances. Moreover, LSPS was optimal with a large battery capacity and close to optimal for a low export rate and peak demand charge. With 0 or low peak demand charges, all control algorithms performed close to optimal. In comparison, the RATP and backup algorithms were sensitive to changes in battery capacity, salvage rate, peak demand charge, and export rate when the building entered the net-producing regime. For cases where the battery capacity was large or the peak price was low, RATP outperformed RL and approached optimal with deterministic predictions. However, with a high salvage value and demand price, the RATP algorithm’s performance degraded significantly and became even worse than the backup mode. Backup mode varied mostly linearly in its gap to the optimal case. It performed well only in special cases with high salvage value rates and low peak demand cases.